The probability that chester will find the hidden treasure is 7/10 what are the odds that cheaster will find the hidden treasure
Answer by stanbon(75887)

The probability that chester will find the hidden treasure is 7/10 what are the odds that cheaster will find the hidden treasure
-------
P(find) = 7/10
P(not find) = 3/10
----
odds to find = P(find)/P(not find) = 7:3
=======================
Cheers,
Stan H.